# Unleashing the Power of Meta-tuning for Few-shot Generalization Through Sparse Interpolated Experts

### v1

@@ 5 Experiment
5.2 Roles of sparsity in meta-tuning with SMAT.
@@
ID vs. OOD tradeoff through controlled sparsity levels.    We observe that adjusting the expert sparsity $\tau$ allows us to balance the trade-off between in-domain (ID) and out-of-domain (OOD) performance of our meta-tuned SMAT models. In Fig. 2(a), we see that the OOD performance generally improves while the ID performance decreases as the expert sparsity level $\tau$ increases. We hypothesize that this result is due to the stronger intrinsic meta-regularization effect associated with higher sparsity constraints, as well as the better preservation of the more generic pre-trained features through weight interpolation between the meta-tuned and pre-trained parameters in our formulation in sparse interpolated experts (see Eqn. (3.1)). Both of these factors help to mitigate meta-overfitting to the ID meta-training tasks hence improve meta-generalization.
@@
Sparsity in experts encourages specialization.    Meta-learned sparsity patterns in the experts $\bm{z}_{m}\odot\bm{\theta}^{\delta}$ induce sparsity in the meta-gradients which alleviates harmful task-interference, thereby encouraging specialization among the experts. In Fig. 2(b), we compare the average alignment of meta-gradients between tasks during training for two SMAT models with different expert sparsity levels $\tau$, where the alignment of gradients, defined as $\mathbb{E}_{{\mathcal{T}}_{i},{\mathcal{T}}_{j}\sim{\mathcal{P}}_{ID}}[\text{cos}(\nabla_{\bm{\theta}}{\mathcal{L}}_{i},\nabla_{\bm{\theta}}{\mathcal{L}}_{j})]$, is respectively computed for $\bm{\theta}\in\{\bm{\theta}^{\delta},\bm{z}_{m}\odot\bm{\theta}^{\delta},\ \forall m\in[|{\mathcal{M}}|]$, i.e., the overall meta-tuned parameter and each expert individually. The results show that the higher sparsity level $\tau=0.9$ in SMAT can lead to greater alignment of meta-gradients between tasks. Moreover, the alignments in the experts’ meta-gradients (which are sparse) are generally higher than that of the overall one, (i.e., w.r.t. $\bm{\theta}^{\delta}$ in black) - a sign for development of each expert into highly specialized parameters.
@@
(a) Average performance tradeoff on sampled ID vs OOD tasks as a function of (color) expert sparsity level $\tau$, and (marker) number of experts.
@@
(b) Meta-gradients alignment between tasks throughout for SMAT with low and high sparsity levels. Meta-gradients are calculated w.r.t. the parameters shown in the legend.
Figure 3: The functions of sparsity constraints on experts in SMAT. (a) The expert sparsity levels establish a trade-off between out-of-distribution performance and in-distribution performance. (b) The sparsity in experts promotes expert specialization as a result of highly aligned meta-gradients w.r.t. the parameters of each expert.
5.3 Ablation studies.
@@
Number of experts in SMAT.    In Fig. 2(a), we see that having more experts, hence higher model capacity given the same expert sparsity $\tau$, generally increases both ID and OOD performance of our model. The aforementioned ID vs. OOD tradeoff still exists for different numbers of experts; however, the OOD-to-ID tradeoff ratio (defined as $\frac{\Delta_{\text{OOD acc}}}{\Delta_{\text{ID acc}}}$) varies - with $|{\mathcal{M}}|=4$ experts having the worst tradeoff ratio, and increasing $|{\mathcal{M}}|$ from $4\rightarrow 8$ leads to the most significant gain in the ratio while the improvement seems to saturate when further increasing $|{\mathcal{M}}|$ from $8\rightarrow 16$.
Scale of the Meta-tuning Datasets.    In Fig. 4, we investigate the impact of the quantity and diversity of tasks observed during meta-tuning on the generalization performance of meta-tuned models on unseen meta-testing tasks. As anticipated, the overall generalization performance (on ID and OOD testing tasks), of both PMF and ours, improves as the scale of the meta-tuning datasets increases along the axis. However, even with increased quantity and diversity in the meta-tuning tasks, the OOD performance of PMF is not always better than that of the pre-trained model, which we conjecture is due to both meta-overfitting and harmful task inference. In constrat, SMAT consistently achieves better OOD, and overall, performance compared to both PMF and pre-trained models, with a noticeable $\sim$4improvement in relative OOD performance even under the low-diversity (IO) scenario. It is worth noting that this is also where the largest improvement in OOD performance for SMAT occurs. Intuitively, this is because Omniglot is very different from ImageNet - the only training source prior to its addition, resulting in a significant increase in task diversity. With only these two datasets, SMAT achieves comparable OOD performance to its full-MD version. These results highlight the effectiveness of using SMAT for meta-tuning under low data diversity settings, as well as its ability to achieve improved few-shot generalization by better leveraging task diversity during meta-tuning.
@@
Figure 5: (a-b) Model capacity (i.e., number of non-zero parameters) grouped by a): layer types, and b): layer depth. (c-d) Expert specialisation. c) Dendrogram of task similarity computed on different tasks. d) Overlap between masks.
@@ 5.4 Qualitative visualization
Patterned sparsity emerges through meta-tuning.    In Fig. 5, we visualize the sparsity patterns on masks identified through meta-tuning. We observe that the sparsity levels vary significantly depending on the layer types (a), and depths (b). Specifically, the intermediate layers (5-9) have lower per-layer sparsity, while the first few layers are highly sparsified, with sparsity levels as high as 95%. Among the different layer types, we find that three types of layers retain most of their modulation parameters (non-zeros): (1) the first input embedding layers, (2) values of the attention module, and (3) linear layers of attention and feedforward modules. Across different masks, we notice that the standard deviations of sparsity levels are particularly larger for the first embedding layers and throughout layers at all depths. By examining the overlapping ratio (defined as $\frac{|(\bm{z}_{i}\cap\bm{z}_{j})\neq 0|}{|(\bm{z}_{i}\cup\bm{z}_{j})\neq 0|}$), as shown in Appendix Fig. 5(d), we find that different masks, hence experts, generally have a small overlap. This indicates that SMAT has indeed discovered a diverse set of sparse interpolated experts through meta-tuning.
@@ 6 Conclusion
7 Broader Impact
@@ A.1 SMAT for meta-training
The pseudocode for meta-tuning using SMAT can be found in Alg. 1 below. Upon acceptance of the paper, we will make our code implementation publicly available..
@@
  Data: Meta-training tasks ${\mathbb{T}}:=\{{\mathcal{T}}_{1},{\mathcal{T}}_{2},...{\mathcal{T}}_{N}\}$
@@
  Output: Variational distribution $\Phi=\{\phi_{m}\}^{M}_{m=1}$ for the sparse masks; Hypernetwork $h_{\psi}$; Meta-tuned parameters $\bm{\theta}^{\delta}$
@@
  Initialize the Lagragian multipliers $\lambda_{m}=0,\ \forall m\in[M]$ for sparsity constraint $L_{0}(z_{m}\sim q_{\phi_{m}})\leq\tau$.
@@
        ${\mathcal{T}}_{i}\rightarrow{\mathcal{T}}^{s}_{i}:=\{{\mathcal{X}}_{i}^{s},{\mathcal{Y}}_{i}^{s}\}^{N_{i}^{s}},{\mathcal{T}}_{i}^{q}:=\{{\mathcal{X}}_{i}^{q},{\mathcal{Y}}_{i}^{q}\}^{N_{i}^{q}}$# Split into support and query sets
@@
     ${\mathcal{C}}o_{m}:=[\frac{\lambda}{\text{\small dim}(\bm{\phi}_{m})}\sum_{j}^{\text{dim}(\bm{\phi}_{m})}(1-Q_{\bm{\phi}_{m}}(s_{j}\leq 0))]-\tau,\ \forall m\in[M]$# sparsity loss for each expert
@@
           $\lambda_{m}\leftarrow\mathtt{GradientAscent}(\nabla_{\lambda_{m}}{\mathcal{C}}o_{m}\frac{1}{|\mathbb{B}|}\sum_{i}\alpha_{i,m})$# Maximization problem in Eqn. (3.1)
@@
            # the constraint $L_{0}(z_{m}\sim q_{\phi_{m}})\geq\tau$ is satisfied
@@ A.2 SMAT for meta-testing
  $*l*$ = positive infinity  # use to record the lowest support loss during exploration
@@ Appendix B Experiment details
B.2 Details for meta-testing.
@@ Appendix C Additional experimental results
C.1 Meta-dataset meta-testing results for meta-tuned supervised pre-trained backbones.
@@
C.2 Different fine-tuning strategies for SMAT.
    In Tab 4 below, we evaluate various meta-testing fine-tuning strategies for SMAT and compare their performance. We first note the effectiveness of our proposed gradient-free expert selection method (see Section 3.2), as evidenced by its improved performance compared to directly using SMAT with ProtoNet (Snell et al., 2017) for meta-testing. Second, using $\theta_{i}$ as the initialization for full fine-tuning, which has the same capacity as the pre-trained model $\theta^{\mathtt{pre}}$, leads to improved performance over fine-tuning the full SMAT model jointly (i.e., (4)), as it is sufficiently expressive while much more parameter-efficient than the latter hence avoids potential over-fitting issues.

### v3

@@ 5.1 SMAT achieves new SOTA performance
5.2 Roles of sparsity in meta-tuning with SMAT
@@
ID vs. OOD tradeoff through controlled sparsity levels.    We observe that adjusting the expert sparsity $\tau$ allows us to balance the trade-off between in-domain (ID) and out-of-domain (OOD) performance of our meta-tuned SMAT models. In Fig. 3, we see that the OOD performance generally improves while the ID performance decreases as the expert sparsity level $\tau$ increases. We hypothesize that this result is due to the stronger intrinsic meta-regularization effect associated with higher sparsity constraints, as well as the better preservation of the more generic pre-trained features through weight interpolation between the meta-tuned and pre-trained parameters in our formulation in sparse interpolated experts (see Eqn. (4.1)). Both of these factors help to mitigate meta-overfitting to the ID meta-training tasks hence improve meta-generalization.
@@
Sparsity in experts encourages specialization.    Meta-learned sparsity patterns in the experts $\bm{z}_{m}\odot\bm{\theta}^{\delta}$ induce sparsity in the meta-gradients which alleviates harmful task-interference, thereby encouraging specialization among the experts. In Fig. 4, we compare the average alignment of meta-gradients between tasks during training for two SMAT models with different expert sparsity levels $\tau$, where the alignment of gradients, defined as $\mathbb{E}_{{\mathcal{T}}_{i},{\mathcal{T}}_{j}\sim{\mathcal{P}}_{ID}}[\text{cos}(\nabla_{\bm{\theta}}{\mathcal{L}}_{i},\nabla_{\bm{\theta}}{\mathcal{L}}_{j})]$, is respectively computed for $\bm{\theta}\in\{\bm{\theta}^{\delta},\bm{z}_{m}\odot\bm{\theta}^{\delta},\ \forall m\in[|{\mathcal{M}}|]$, i.e., the overall meta-tuned parameter and each expert individually. The results show that the higher sparsity level $\tau=0.9$ in SMAT can lead to greater alignment of meta-gradients between tasks. Moreover, the alignments in the experts’ meta-gradients (which are sparse) are generally higher than that of the overall one, (i.e., w.r.t. $\bm{\theta}^{\delta}$ in black) - a sign for development of each expert into highly specialized parameters.
Table 2: Few-shot testing results on the Meta-dataset benchmark for various SMAT meta-tuned Vision Transformer backbones.
with gradient-based full fine-tuning
SMAT meta-tuned ViT backbone Params. In-domain Out-of-domain Avg
INet Acraft Omglot CUB DTD Qdraw Fungi Flower Sign Coco Cifar10 Cifar100 Mnist
DINO Small (Caron et al., 2021) 21M 75.24 90.01 90.83 92.57 88.41 79.17 73.31 97.21 91.33 63.11 92.08 85.91 96.73 86.42
DINO Base (Caron et al., 2021) 86M 78.28 92.19 92.89 93.41 86.97 80.35 75.59 96.93 91.23 64.85 92.57 86.95 97.70 87.22
Sup21k Small (Dosovitskiy et al., 2020) 21M 82.10 89.16 88.56 94.98 89.05 79.36 74.50 99.19 90.13 68.19 94.87 90.32 96.76 87.81
Sup21k Base (Dosovitskiy et al., 2020) 86M 85.01 87.41 88.05 95.52 87.59 79.78 74.01 99.34 91.38 68.51 93.61 92.84 96.88 87.96
Sup21k Large (Dosovitskiy et al., 2020) 307M 87.38 88.00 88.16 95.59 88.56 81.02 74.34 99.09 91.77 69.50 95.27 92.62 96.79 88.43
@@
Figure 3: Average performance tradeoff on sampled ID vs OOD tasks as a function of (color) expert sparsity level $\tau$, and (marker) number of experts.
@@
Figure 4: Meta-gradients alignment between tasks throughout for SMAT with low and high sparsity levels. Meta-gradients are calculated w.r.t. the parameters shown in the legend.
Figure 5: (a-b) Meta-learned model capacity after meta-tuning (i.e., number of non-zero parameters) grouped by a): layer types, and b): layer depth. (c-d) Expert specialisation. c) Dendrogram of task similarity based on expert selection scores. d) Overlap between masks.
5.3 Ablation studies
@@
Number of experts in SMAT.    In Fig. 3, we see that having more experts, hence higher model capacity given the same expert sparsity $\tau$, generally increases both ID and OOD performance of our model. The aforementioned ID vs. OOD tradeoff still exists for different numbers of experts; however, the OOD-to-ID tradeoff ratio (defined as $\frac{\Delta_{\text{OOD acc}}}{\Delta_{\text{ID acc}}}$) varies - with $|{\mathcal{M}}|=4$ experts having the worst tradeoff ratio, and increasing $|{\mathcal{M}}|$ from $4\rightarrow 8$ leads to the most significant gain in the ratio while the improvement seems to saturate when further increasing $|{\mathcal{M}}|$ from $8\rightarrow 16$.
Scale of Meta-tuning datasets.    In Fig. 6, we investigate the impact of the quantity and diversity of tasks observed during meta-tuning on the generalization performance of meta-tuned models on unseen meta-testing tasks. As anticipated, the overall generalization performance (on ID and OOD testing tasks), of both PMF and our model, improves as the scale of the meta-tuning datasets increases along the x-axis. However, even with increased quantity and diversity in the meta-tuning tasks, the OOD performance of PMF is not always better than that of the pre-trained model, which we conjecture is due to both meta-overfitting and harmful task interference.
@@
In contrast, SMAT consistently achieves better OOD and overall performance compared to both PMF and pre-trained models, with a noticeable $\sim$4improvement in relative OOD performance even in the low-diversity (IO) scenario. It is worth noting that this is also where the largest improvement in OOD performance for SMAT occurs. Intuitively, this is because Omniglot is very different from ImageNet, which was the only training source prior to its addition, resulting in a significant increase in task diversity. With only these two datasets, SMAT achieves comparable OOD performance to its full-MD version. These results highlight the effectiveness of using SMAT for meta-tuning in low data diversity settings, as well as its ability to achieve improved few-shot generalization by better leveraging task diversity during meta-tuning.
Scale of Vision Transformer backbones.  In Tab. 2, we present the meta-testing results on MD of various Vision Transformer backbones meta-tuned with SMAT. Our findings indicate that the larger models tend to offer superior overall performance but may require substantially more computational resources due to the number of parameters. These results can guide the selection of appropriate transformer architectures based on specific application requirements and resource constraints.
@@ 5.4 Qualitative visualization
Patterned sparsity emerges through meta-tuning.    In Fig. 5, we visualize the sparsity patterns on masks identified through meta-tuning. We observe that the sparsity levels vary significantly depending on the layer types (a), and depths (b). Specifically, the intermediate layers (5-9) have lower per-layer sparsity, while the first few layers are highly sparsified, with sparsity levels as high as 95%. Among the different layer types, we find that three types of layers retain most of their modulation parameters (non-zeros): (1) the first input embedding layers, (2) values of the attention module, and (3) linear layers of attention and feedforward modules. Across different masks, we notice that the standard deviations of sparsity levels are particularly larger for the first embedding layers and throughout layers at all depths. By examining the overlapping ratio (defined as $\frac{|(\bm{z}_{i}\cap\bm{z}_{j})\neq 0|}{|(\bm{z}_{i}\cup\bm{z}_{j})\neq 0|}$), as shown in Fig. 5(d), we find that different masks, hence experts, generally have a small overlap. This indicates that SMAT has indeed discovered a diverse set of sparse interpolated experts through meta-tuning.
@@ 6 Conclusion
Acknowledgement
This work was supported by the Research Matching Grant Scheme (RMGS 9229111) founded by the University Grants Committee of Hong Kong, and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2022-0-00713, Meta-learning applicable to real-world problems)
Impact Statement
@@ A.1 SMAT for meta-training
The pseudocode for meta-tuning using SMAT can be found in Alg. 1 below. Our implementation is publicly available at github.com/szc12153/sparse_meta_tuning.
@@
  Data: Meta-training tasks ${\mathbb{T}}:=\{{\mathcal{T}}_{1},{\mathcal{T}}_{2},...{\mathcal{T}}_{N}\}\sim\mathcal{P}_{ID}$
@@
  Output: Variational distribution $\Phi=\{\phi_{m}\}^{|{\mathcal{M}}|}_{m=1}$ for the sparse masks; Hypernetwork $h_{\psi}$; Meta-tuned parameters $\bm{\theta}^{\delta}$
@@
  Initialize the Lagrangian multipliers $\lambda_{m}=0$ for sparsity constraint $1-\frac{L_{0}(\bm{z}_{m})}{\text{dim}(\bm{z}_{m})}\geq\tau,\ \ \forall m\in[|{\mathcal{M}}|]$.
@@
        ${\mathcal{T}}_{i}\rightarrow{\mathcal{T}}^{s}_{i},{\mathcal{T}}_{i}^{q}$# Split into support and query sets
@@
     ${\mathcal{C}}o_{m}:=\frac{\lambda}{\text{\small dim}(\bm{\phi}_{m})}\sum_{j}^{\text{dim}(\bm{\phi}_{m})}Q_{\bm{\phi}_{m}}(s_{j}\leq 0)-\tau,\ \forall m\in[M]$# sparsity loss for each expert
@@
           $\lambda_{m}\leftarrow\mathtt{GradientAscent}(\nabla_{\lambda_{m}}{\mathcal{C}}o_{m}\frac{\sum_{i\in\mathbb{B}}\alpha_{i,m}}{|\mathbb{B}|})$# Maximization problem in Eqn. (4.1)
@@
            # expected expert sparsity is at least $\tau$, the constraint is satisfied
@@ A.2 SMAT for meta-testing
  $l*$ = positive infinity  # use to record the lowest support loss during exploration
@@ B.1 Implementation of baselines
B.2 Details for meta-testing
@@ Appendix C Additional experimental results
C.1 Meta-dataset meta-testing results for meta-tuned supervised pre-trained backbones
@@
C.2 Different fine-tuning strategies for SMAT at meta-testing
Performance.    In Tab 5 below, we evaluate various meta-testing fine-tuning strategies for SMAT and compare their performance. We first note the effectiveness of our proposed gradient-free expert selection method (see Section 4.2), as evidenced by its improved performance compared to directly using SMAT with ProtoNet (Snell et al., 2017) for meta-testing. Second, using $\theta_{i}$ as the initialization for full fine-tuning, which has the same capacity as the pre-trained model $\theta^{\mathtt{pre}}$, leads to improved performance over fine-tuning the full SMAT model jointly (i.e., (4)), as it is sufficiently expressive while much more parameter-efficient than the latter hence avoids potential over-fitting issues. We provide more explaination as follows.
Gradient-based fine-tuning outperforms gradient-free fine-tuning primarily because it is much more flexible. The extra flexibility of gradient-based fine-tuning (Alg. 3, Appendix) stems from the fact that the entire model $\theta_{i}$ is allowed to be updated, whereas gradient-free fine-tuning (Alg. 2, Appendix) only allows updates on the expert selection weights $\alpha_{i}$ with dimensions equal to the number of experts in SMAT; while all other parameters remain frozen. That said, the hypothesis space of our gradient-free fine-tuning algorithm is much more tightly constrained around the model meta-tuned on ID tasks. As a result, the effectiveness of gradient-free fine-tuning is limited on certain OOD tasks that exhibit noticeable distribution shift, such as TrafficSign. In such cases, the meta-tuned ID model may become inadequate, requiring significant parameter updates.
@@ C.2 Different fine-tuning strategies for SMAT at meta-testing
Computational cost.     We present results below in Tab. 6 for a quantitative comparison of computational cost, in terms of time and GPU memory, between w/o fine-tuning, gradient-free fine-tuning, and gradient-based full fine-tuning for SMAT. All fine-tuning are carried out in FP16 mixed precision. In particular, our gradient-free fine-tuning method offers a $+1.59$% improvement on average over w/o fine-tuning at no additional memory cost, while saves $3/4$ of the total memory cost of gradient-based fine-tuning. Gradient-based fine-tuning, however, outperforms both by at least $+10.31$% despite requiring 4x the GPU memory of both, and 2x the time cost compared to gradient-free fine-tuning.
Table 6: Computational cost for fine-tuning a SMAT model (ViT-DINO-Small backbone) at meta-testing time with different fine-tuning strategies.
Method Time (sec./task) GPU memory (MiB) Avg. Acc (Tab. 5)
w/o fine-tuning (ProtoNet) 0.2 4332 73.83
Alg.( 2). gradient-free fine-tuning (50steps) 6.4 4332 75.43
Alg.( 3). gradient-based full fine-tuning (50steps) 11.8 17264 85.74
@@
Appendix D Sparse interpolated experts
As previously stated, by rearranging Eqn. (1) slightly, the experts now essentially become different sparse interpolations between the same pre-trained and meta-tuned models. Here are the details.
Starting from Eqn. (1), we have:
$\bm{\theta}_{i}$ $=\bm{\theta}^{\mathtt{pre}}+\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}\bm{\theta}^{\delta}_{m}$ (3)
$=\bm{\theta}^{\mathtt{pre}}+\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{z}_{m}\odot\bm{\theta}^{\delta})$ (4)
$=\bm{\theta}^{\mathtt{pre}}+\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{z}_{m}\odot\bm{\theta}^{\delta+\mathtt{pre}})-\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{z}_{m}\odot\bm{\theta}^{\mathtt{pre}})$ (5)
$=\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{1}\odot\bm{\theta}^{\mathtt{pre}})+\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{z}_{m}\odot\bm{\theta}^{\delta+\mathtt{pre}})-\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}(\bm{z}_{m}\odot\bm{\theta}^{\mathtt{pre}})$ (6)
$=\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}((\bm{1}-\bm{z}_{m})\odot\bm{\theta}^{\mathtt{pre}}+\bm{z}_{m}\odot\bm{\theta}^{\delta+\mathtt{pre}}).$ (7)
[…]